HOME
ESR
EXPERIMENTS
LABS etc
|
E121
E127
E132
E125
E143
laser_cooling_2021
E142
Ê128
E146
E0052
E0018
E0028
E0038
E0075
E0203_E0018
|
masses & isomeres 2025, Page 5 of 8 |
Not logged in |
 |
|
New entries since: | Thu Jan 1 01:00:00 1970 | |
Wed Apr 2 13:35:17 2025, Michael, Alex, Maeve, DAQ, RSAs data taking
|
Wed Apr 2 13:27:52 2025, Michael, Alex, Maeve, Accelerator, No beam in the ESR
|
Wed Apr 2 12:58:15 2025, Ruijiu, Wolfram, Analysis, frequency drift  
|
Wed Apr 2 12:13:57 2025, Michael, Alex, Accelerator, Some empty injections triggered while beam patterns are changed
|
Wed Apr 2 11:09:42 2025, Michael, Alex, Accelerator, The beam is back in the ESR
|
Wed Apr 2 10:55:16 2025, Michael, Alex, Wolfram, Accelerator, No beam in the ESR
|
Wed Apr 2 10:31:15 2025, Michael, Alex, Accelerator, The beam is back from 10:10
|
Wed Apr 2 10:05:06 2025, Michael, Alex, Yuri, Accelerator, No beam since 09:59
|
Wed Apr 2 10:02:40 2025, Sergey + Yuri, Accelerator, ESR Timeout
|
Wed Apr 2 09:28:59 2025, Michael, Alex, General, Empty injections in the ESR
|
Wed Apr 2 08:47:27 2025, Michael, Alex, General, The beam is back
|
Wed Apr 2 08:04:47 2025, Michael, Alex, General, No beam
|
Wed Apr 2 07:41:10 2025, Michael, Alex, General, Frequency shift in Schottky
|
Wed Apr 2 07:39:41 2025, Michael, Alex, General, The beam is back at 07:35
|
Wed Apr 2 07:33:59 2025, Michael, Alex, General, Beam gone again
|
Wed Apr 2 07:25:15 2025, Carlo, Chris, Alex, Michael , General, Beam back
|
Wed Apr 2 07:16:33 2025, Carlo, Chris, Alex, Michael, DAQ, ESR not requesting
|
Wed Apr 2 05:48:31 2025, Chris, Carlo, Zac, General, Beam returned
|
Wed Apr 2 04:59:50 2025, Zac, Chris, Carlo, General, Receiving injections but not seeing anything on RSAs 
|
Wed Apr 2 04:33:57 2025, Carlo, Zac, Chris, DAQ, Screenshots Lassie 
|
|